Population studies of fiddler crabs on a tropical island reported 1.1 x 10^4 in 1980 and 1.5 x 10^4 in 1982. Predict the time that the maximum population, 2.0 x 10^4, that can be supported by the island's resources will be reached.
Answer by stanbon(48569)

Population studies of fiddler crabs on a tropical island reported 1.1 x 10^4 in 1980 and 1.5 x 10^4 in 1982. Predict the time that the maximum population, 2.0 x 10^4, that can be supported by the island's resources will be reached.
-------------------------
Generate an equation based on (1982, 1.5x10^4) and (1980, 1.1x10^4)
-----------------
slope = (1.5x10^4 - 1.1x10^4) / (1982-1980) = 0.4x10^4/2 = 2000
-----------------------------------
So the crab population is growing by 2000 each year.
-----------------
2.0x10^4 - 1.5x10^4 = 0.5(1x10^4) = (1/2)(10,000 = 5000
-----------------------
5000/(2000 per year) = 2.5 years
=================================
The population will be 2x10^4 in 1982 + 2.5 = 1984.5 or 1985
=================================================================
Cheers,
Stan H.
========